# Improving Grip Stability Using Passive Compliant Microspine Arrays for Soft Robots in Unstructured Terrain

### v1

@@ Abstract
Microspine grippers are small spines commonly found on insect legs that reinforce surface interaction by engaging with asperities to increase shear force and traction. An array of such microspines, when integrated into the limbs or undercarriage of a robot, can provide the ability to maneuver uneven terrains, traverse inclines, and even climb walls. Conformability and adaptability of soft robots makes them ideal candidates for these applications involving traversal of complex, unstructured terrains. However, there remains a real-life realization gap for soft locomotors pertaining to their transition from controlled lab environment to the field by improving grip stability through effective integration of microspines. We propose a passive, compliant microspine stacked array design to enhance the locomotion capabilities of mobile soft robots, in our case, ones that are motor tendon actuated. We offer a standardized microspine array integration method with effective soft-compliant stiffness integration, and reduced complexity resulting from a single actuator passively controlling them. The presented design utilizes a two-row, stacked microspine array configuration that offers additional gripping capabilities on extremely steep/irregular surfaces from the top row while not hindering the effectiveness of the more frequently active bottom row. We explore different configurations of the microspine array to account for changing surface topologies and enable independent, adaptable gripping of asperities per microspine. Field test experiments are conducted on various rough surfaces including concrete, brick, compact sand, and tree roots with three robots consisting of a baseline without microspines compared against two robots with different combinations of microspine arrays. Tracking results indicate that the inclusion of microspine arrays increases planar displacement on average by 15 and 8 times over the baseline with the two microspine designs respectively, improves locomotion repeatability, and, critically, consistently increases terrain traversability.
@@ I INTRODUCTION
In nature, animals resist slipping and falling by increasing interaction with surfaces in a number of ways. Snakes maintain a large surface area in contact with the ground to increase the amount of propulsive force aiding in locomotion. Snake inspired robots are developed with textured skins for increased traction [1, 2]. However, the efficacy is unknown for limbed robots equipped with these skins and will require additional coordination. Geckos utilize van der Waals interactions to provide an adhesive force against surfaces. Attaching gecko inspired, synthetic adhesives to extremities of robots is an attractive option due to the range of surfaces the robots can adhere to [3, 4, 5, 6, 7]. These directional adhesion mechanisms are sensitive to wear and tear and require periodic cleaning for reliable adhesion. Insects such as caterpillars and cockroaches are able to climb vertical surfaces with small spines attached to their legs. These microspine grippers increase shear force and traction by engaging with surface asperities. They do not penetrate surfaces, but rather reinforce surface interaction by gripping onto jagged, microscopic edges. Biological studies have looked at the relationship between microspine arrays and surface interaction, comparing the efficacy on different smooth and rough surfaces [8]. The outcome from these studies encourage the creation of synthetic microspine arrays to enable maneuverability of uneven terrains, traversability of inclines, and climbing walls [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. These arrays can become very complicated, and there exists a trade-off between design complexity and surface adaptability for many of these robots as shown in Fig. 1.
Climbing and crawling rigid robots developed in the past decade employ similar methods to combat slipping along vertical walls [18, 20, 21, 22, 23]. JPL’s LEMUR IIB, a four-limbed robot, features multiple grippers, each equipped with at least 250 microspines capable of adhering to both convex and concave surfaces. Each microspine is suspended with a steel hook, allowing it to stretch and move relative to its neighbors to find suitable surfaces to grip. This enables the robot to climb vertically and traverse on cave ceilings. The shear forces from the microspines integrated in the gripper gives the ability to support heavy loads [24]. These grippers possess the necessary traction to navigate microgravity environments, such as asteroids, comets, and small moons, where conventional ground pressure is lacking [25]. Microspines latch into small asperities on a rock’s surface such as pits, cracks, slopes, or any kind of topography that will snag a hook. Success of an individual spine catching is dependent on surface roughness, surface geometry, and asperity distribution [26].
Integrating microspines into Soft Robots (SoRos) is an attractive option due to their adaptability and conformability to changing surface topologies, highlighted in Fig. 2. The continuum nature and impact resistance of soft materials passively allow SoRos additional flexibility and more effective interaction with complex and non-uniform surfaces. However, SoRos lack grip stability, contributing to them historically struggling with efficient locomotion as well as locomoting over unstructured terrain. Because of this, SoRo designs that can traverse outside and perform real tasks outside of a lab setting are under-researched. Utilizing microspines has the potential to improve traction, increase grip stability, and provide the ability to reliably maneuver real-world terrains. One of the main design challenges pertains to integrating a soft, low stiffness body with hard, high stiffness micropsines. A two segment, wriggling SoRo avoids this challenge by adhering an array of dual material, “soft microspines” made of rubber along the ventral of the body to successfully increase anisotropic friction [27]. A starfish inspired SoRo implements a similar technique by including soft, tube feet reminiscent of microspines along the entire underside of the five-limbed robot. Magnetization of the tube feet allows omnidirectional movement and reduces the motion resistance from the ground, enhancing the adaptability of the SoRo on different surfaces [28]. Embedding hard objects in soft materials through intelligent mechanical design is necessary to take advantage of the benefits of hard spines without hindering the soft deformable properties. A soft inchworm design attaches an array of microspines to either foot of the inchworm using adhesive bonding technology [29]. However, deeply irregular surfaces remain difficult if not impossible to overcome due to the uniform distribution of the microspines and integration technique that restricts the usage to surfaces with regular, fine asperities. This shows the need for compliance and independent movement per microspine to increase surface geometry traversability.
@@
Figure 1: Surface adaptability vs. design complexity with various microspine robots [9, 13, 25, 19, 29, 30, 31] compared against one of the presented designs, 1ML.
@@
Contributions. We aim to bridge the realization gap by attaching microspine technology onto the tips of soft Motor Tendon Actuated (MTA) limbs, vastly improving the grip stability and types of traversable terrain of mobile SoRos. We propose an elegant, single-material design with intelligent soft-compliant integration that reduces complexity by using a single actuator to passively control an entire array fixed to the end of a soft limb. In this research, we (1) propose a compliant mechanism, two-row stacked microspine array design that improves grip stability and increases traversable surface topologies of mobile SoRos; (2) identify critical design parameters that improve locomotion capabilities while reducing complexity by controlling an entire microspine array with only a single actuator through intelligent soft-compliant integration; (3) investigate the grip stability and repeatability of a baseline SoRo compared against two different microspine array configurations on uniform concrete, partially uniform brick, granular compact sand, and non-uniform tree roots; and (4) analyze tracking results that indicate the inclusion of compliant microspine arrays in SoRos increases planar displacement on all surfaces through enhanced surface engagement resulting in capabilities to traverse complex, unstructured environments.
@@ II MICROSPINE ARRAY AND ROBOT DESIGNS
There are several design parameters that impact the effectiveness of microspines. The critical ones include (1) adding compliance to individual microspines and the angle at which the microspines interact with a surface, (2) the array configuration and (3) effective integration with the robot body. Even with an optimal design, the microspine array will not be effective on every surface. The aspects that are out of the hands of the designer, but also highly impact surface engagement, include surface roughness, distribution of asperities, and size of asperities.
@@ II-A Compliant Mechanism Microspine Design
The single-material mechanism, shown in Fig. 3, allows compliance with an exposed joint while simplifying the fabrication process over previous microspine designs. The compliant mechanism is fabricated with an FDM 3D printer and TPU with 95A Shore hardness. Halfway through the additive manufacturing process, the print is paused. The microspine is inserted into a channel left in the middle of the mechanism, highlighted in Fig. 3c), and the print is resumed. Once finished, the angle of the bare microspine can be modified for different surface topologies while the body remains secure in the mechanism. The angle of surface interaction, $\alpha$, was fixed at roughly $45^{\circ}$ during testing.
Figure 3: Compliant mechanism design. a) A hinge joint enables passive compliance. b) Holes embedded on the righthand side of the mechanism allow anchoring into the silicone limb. c) A microspine is inserted in a center channel matching the spine topology set halfway into the mechanism.
@@ II-B Microspine Array Configuration
The array configuration ensures multiple microspines remain active on various complex surfaces. We propose a two-row, stacked array configuration consisting of ten microspines with four on the top row and six on the bottom. The microspines on the bottom row are commonly active on more uniform terrain. The top row can become active on steep/highly irregular surfaces without hindering the movement of the bottom row of microspines. The critical parameter when designing the array configuration is ensuring adequate surface interaction and gripping regardless of topology. Crucially, not all microspines need to interact with a surface for the microspine array to be effective, shown in Fig. 4. This is a byproduct of the passive nature and built-in redundancy of the system.
Figure 4: Two-row, stacked array configuration. a) Close-up of the microspines gripping onto a non-uniform rock. b) This diagram highlights which microspines are interacting with the surface. On this rock, all 6 of the microspines on the bottom row are active. c) Close-up of the microspines gripping onto a steeper rock. d) On this rock, 2 of the bottom row and 3 of the top row microspines are active.
@@ II-C Effective Soft-Compliant Integration Through Anchoring
The soft-compliant integration reduces design complexity by allowing each microspine to passively move independent of one another with a single actuator controlling the entire array configuration. To achieve this, a mold is created with channels for each microspine compliant mechanism to attach to the tip of a SoRo limb. The SoRo prototype used for experimentation is cast out of DragonSkin™  silicone rubber using a custom mold, shown in Fig. 5. Therefore, a modified limb mold is used for integrating the microspine array in a consistent, standardized manner. Half of the compliant mechanism contains holes that mechanically anchor it into the silicone limb, highlighted in Fig. 3b), preventing it from being freely pulled out of the limb during microspine gripping. This anchoring method is essential for ensuring the microspine does not come loose over time. The remaining, exposed half of the mechanism contains the microspine.
@@ II-D Soft Robot Design
A tetherless, three limb MTA SoRo with on-board power and processing with AprilTags on each limb is used as the experimental prototype. The components of the physical robot are shown in Fig. 6. Outward trapezoid cavities are introduced on the underside of each limb to provide optimal stiffness and curling ability. This allows the robot to lift the limb and electronic payload. The use of MTA for body deformation enables reliable and efficient limb actuation. The reader may refer to [32] for more details.
@@ III EXPERIMENTATION
III-A Experimental Prototypes
The baseline experimental prototype with zero microspine limbs, 0ML, is a three-limb MTA SoRo coming from a previous work [33]. We compare the baseline against the addition of two different microspine configurations. The first SoRo equipped with a microspine array, 1ML, has them affixed to one limb that acts as the leader. The last SoRo, 2ML, contains microspine arrays equipped on two limbs with these acting as dual leaders. In all microspine configurations, the microspines face opposite the intended direction of movement. CAD models of the three prototypes are shown in Fig. 7.
@@ III-A Experimental Prototypes
A push-pull translation gait, Fig.  8, is used for the experiments. Here the gait sequence involves actuation of one limb followed by actuation of the two relaxed limbs. It is worth reminding the reader that this gait is not optimal for all four surfaces and three prototypes. In a previous work, translation and rotation gaits were discovered for four limb SoRos using an environment-centric framework [33]. The same methodology was used to generate an optimal translation gait for the baseline three limb SoRo (0ML) on a rubber mat. This gait was used on all surfaces for 0ML, 1ML, and 2ML for a fair comparison, but it is expected that much better performance is possible with an optimized gait found for each prototype and surface combination.
Figure 8: Translation gait: one limb (in maroon) actuates while the other two relax, then the previously relaxed limbs actuate while the other relaxes. This is an optimal, 1.1s long gait for the baseline SoRo on a rubber mat [33].
III-B Experimental Setup
The experiments are carried out with an overhead camera attached to a tripod next to the test area, shown in Fig. 9. Field tests are performed on four different surfaces. The starting pose, position and orientation, is same for each SoRo to ensure consistency. Tracking is performed with an AprilTag fixed to the end of each limb serving as fiducial markers. The average displacement per gait as well as overall displacement for each trial run is recorded.
Figure 9: Example of experimental setup for outside field locations in a forest with a dirt patch. The overhead camera is orthogonal to the ground and approximately 4’ above the surface. Experiments took place with the three prototypes discussed previously: 0ML, 1ML and 2ML.
III-C Field Experiments
The field experiments are tested on four rough surfaces around the University of Alabama campus. We look at uniform concrete, partially uniform brick, granular compact sand with pebbles, and a non-uniform forest floor containing leaf litter and large tree roots. Tests are conducted using the push-pull translation gait shown in Fig. 8. Three trials (60 gaits per trial) are performed for a particular configuration and surface with three prototypes and four surfaces, resulting in a total of 36 trials.
A 36h11 family AprilTag from the AprilTag visual fiducial system [34] is attached to each limb where they are readily visible while not hindering the movement. To compensate for occlusion and easy access to the hub, three tags were used to find the robot’s position. Each limb of a robot has a different tag attached which represents a number from $0-8$. Before starting the tracking, the threshold parameters for different background environments are obtained. The tags are detected using an AprilTag detector library in Python. The tracking algorithm and data processing code is available at github.com/AgileRoboticsLab/SoftRobotics-Microspines.

### v2

@@ Abstract
Microspine grippers are small spines commonly found on insect legs that reinforce surface interaction by engaging with asperities to increase shear force and traction. An array of such microspines, when integrated into the limbs or undercarriage of a robot, can provide the ability to maneuver uneven terrains, traverse inclines, and even climb walls. Meanwhile, the conformability and adaptability of soft robots makes them ideal candidates for applications involving traversal of complex, unstructured terrains. However, there remains a real-life realization gap for soft locomotors pertaining to their transition from controlled lab environment to the field that can be bridged by improving grip stability through effective integration of microspines. In this research, a passive, compliant microspine stacked array design is proposed to enhance the locomotion capabilities of mobile soft robots. A microspine array integration method effectively addresses the stiffness mismatch between soft, compliant, and rigid components. Additionally, a reduction in complexity results from actuation of the surface-conformable soft limb using a single actuator. The two-row, stacked microspine array configuration offers improved gripping capabilities on steep and irregular surfaces. This design is incorporated into three different robot configurations - the baseline without microspines and two others with different combinations of microspine arrays. Field experiments are conducted on surfaces of varying surface roughness and non-uniformity - concrete, brick, compact sand, and tree roots. Experimental results demonstrate that the inclusion of microspine arrays increases planar displacement an average of 10 times. The improved grip stability, repeatability, and, terrain traversability is reflected by a decrease in the relative standard deviation of the locomotion gaits.
@@ I INTRODUCTION
In nature, animals resist slipping and falling by increasing interaction with surfaces in a number of ways. Snakes maintain a large surface area in contact with the ground to increase the amount of propulsive force aiding in locomotion. Snake inspired robots are developed with textured skins for increased traction [1, 2]. However, the efficacy is unknown for limbed robots equipped with such skins given the directional nature of the robot-environment interaction resulting from the designs. In constrast, geckos utilize van der Waals interactions to provide an adhesive force against surfaces. Consequently, attaching gecko inspired, synthetic adhesives to extremities of robots has been investigated on range of surfaces [3, 4, 5, 6, 7]. However, these directional adhesion mechanisms are sensitive to wear and tear and require periodic cleaning for reliable adhesion. On the other hand, insects such as caterpillars and cockroaches are able to climb vertical surfaces with small spines attached to their legs. These microspine grippers increase shear force and traction by engaging with surface asperities. They do not penetrate surfaces, but rather reinforce surface interaction by gripping onto jagged, microscopic edges. Biological studies have looked at the relationship between microspine arrays and surface interaction, comparing the efficacy on different smooth and rough surfaces [8]. The outcome from these studies encourage the design of synthetic microspine arrays to enable maneuverability of uneven terrains, traversability of inclines, and climbing walls [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Despite the promise, design and control of these arrays can become very complicated, and there exists a trade-off between design complexity and surface adaptability for many of these robots as shown in Fig. 1.
@@
Figure 1: Surface adaptability vs. design complexity with various microspine robots [9, 13, 20, 19, 21, 22, 23] compared against one of the presented designs in this work, 1ML.
The multimedia attachment can be accessed at https://youtu.be/U9QPgaXc8IU?si=zxycAdzoQ9YBXSA3.
Climbing and crawling rigid robots developed in the past decade employ similar ‘microspine’ methods to combat slipping along vertical walls [18, 24, 25, 26, 27]. For example, JPL’s LEMUR IIB, a four-limbed robot, features multiple grippers, each equipped with at least 250 microspines capable of adhering to both convex and concave surfaces. Here, each microspine is suspended with a steel hook, allowing it to stretch and move relative to its neighbors to find suitable surfaces to grip. Conceptually, the shear forces from the microspines integrated in the gripper gives the ability to support heavy loads [28] and the necessary traction to even navigate microgravity environments [20]. Mechanically, the microspines latch into small asperities on a rock’s surface such as pits, cracks, slopes, or any kind of topography that will snag a hook. Consequently, the success of an individual spine catching is dependent on surface roughness, surface geometry, and asperity distribution [29].
Integrating microspines into Soft Robots (SoRos) is an attractive option due to their adaptability and conformability to changing surface topologies, highlighted in Fig. 2. The continuum nature and impact resistance of soft materials passively allow SoRos additional flexibility and more effective interaction with complex and non-uniform surfaces. However, SoRos lack grip stability, contributing to them historically struggling with efficient locomotion as well as locomoting over unstructured terrain. Because of this, SoRo designs that can traverse outside and perform real tasks outside of a lab setting are under-researched. Integrating microspines has the potential to improve traction, increase grip stability, and provide the ability to reliably maneuver real-world terrains. One of the main design challenges pertains to integrating a soft, low stiffness body with hard, high stiffness micropsines. “Soft microspines” made of rubber along the ventral side of the body increase anisotropic friction and adhere using an array of dual material [30]. Similarily, tube feet reminiscent of microspines along the entire underside of the five-limbed star-fish inspired soft robot use similar technique. Locomotion adaptability can be improved by magnetizing the soft tube feet [31]. Embedding hard objects in soft materials through intelligent mechanical design is critical to take advantage of the benefits of hard spines without hindering the soft deformable properties. A soft inchworm design attaches an array of microspines to either foot of the inchworm using adhesive bonding technology [21]. However, deeply irregular surfaces remain difficult if not impossible to overcome due to the uniform distribution of the microspines and integration technique that restricts the usage to surfaces with regular, fine asperities. All this calls for need of compliance and independent movement per microspine to increase surface geometry traversability.
@@
Contributions. The research addresses the real-world implementation gap by incorporating compliant microspines into the limbs of the Motor Tendon Actuated (MTA) SoRo, significantly enhancing grip stability and traversability across various terrains for mobile SoRos. The proposed single-material design with intelligent soft-compliant integration reduces complexity by using a single actuator to passively control an entire array fixed to the end of a soft limb. This research, (1) proposes a compliant mechanism, two-row stacked microspine array design that improves grip stability and increases traversable surface topologies of mobile SoRos; (2) identifies critical design parameters that improve locomotion capabilities while reducing complexity by controlling an entire microspine array with only a single actuator through intelligent soft-compliant integration; (3) investigates the grip stability and repeatability of a baseline SoRo compared against two different microspine array configurations on uniform concrete, partially uniform brick, granular compact sand, and non-uniform tree roots; and (4) experimentally concludes that the inclusion of compliant microspine arrays in SoRos increases planar displacement on all surfaces through enhanced surface engagement resulting in capabilities to traverse complex, unstructured environments.
@@ II MICROSPINE ARRAY AND ROBOT DESIGNS
Mechanically, the critical design parameters of the mechanism that impact the effectiveness of interaction of the microspines with the environmentcan be identified as (1) compliance of individual microspines and the angle of their interaction with a surface, (2) the array configuration involving multiple microspines, and (3) effective integration with the robot body to manage the rigid-soft stiffness mismatch. The other parameters that are out of the hands of the designer, but also highly impact surface engagement, include surface roughness, distribution of asperities, and size of asperities.
@@ II-A Compliant Mechanism Microspine Design
The single-material mechanism, shown in Fig. 3, allows compliance with an exposed joint while simplifying the fabrication process over others presented in the literature. The compliant mechanism is fabricated with an FDM 3D printer and TPU with 95A Shore hardness. Halfway through the additive manufacturing process, the print is paused. The microspine is inserted into a channel left in the middle of the mechanism, highlighted in Fig. 3c), and the print is resumed. Once finished, the angle of the bare microspine can be modified for different surface topologies while the body remains secure in the mechanism. The angle of surface interaction, $\alpha$, was fixed at roughly $45^{\circ}$ during testing.
Figure 3: Compliant microspine mechanism design. a) A hinge joint enables passive compliance. b) Holes embedded on the righthand side facilitate mechanical integration into the soft limb. c) A rigid microspine is inserted in a center channel matching the spine topology set halfway into the mechanism with contact angle of $\alpha=45\degree$
@@ II-B Microspine Array Configuration
The array configuration ensures multiple microspines remain active on various complex surfaces. The proposed design comprises of two-row, stacked array configuration consisting of ten microspines with four on the top row and six on the bottom. Experimentally, it has been observed that the microspines on the bottom row are commonly active on more uniform terrain. The top row can become active on steep/highly irregular surfaces without hindering the movement of the bottom row of microspines. The critical parameter when designing the array configuration is ensuring adequate surface interaction and gripping regardless of topology. Crucially, not all microspines need to interact with a surface for the microspine array to be effective, shown in Fig. 4. This is a byproduct of the passive nature and built-in redundancy of the system.
Figure 4: Two-row, stacked array configuration. a) Close-up of the microspines gripping onto a non-uniform rock. b) where all six of the microspines on the bottom row are engaged (green) while the top row remains disengaged (red). c) Close-up of the microspines gripping onto a steeper rock. d) where two of the bottom row and three of the top row microspines are engaged with the surface.
@@ II-C Effective Soft-Compliant Integration Through Anchoring
The soft-compliant integration reduces design complexity by allowing each microspine to passively move independent of one another with a single actuator controlling the entire array configuration, equivalently the limb. To achieve this, a mold is created with channels for each microspine compliant mechanism to attach to the tip of a SoRo limb. The SoRo prototype used for experimentation is cast out of DragonSkin™  silicone rubber using a custom mold, shown in Fig. 5. Therefore, a modified limb mold is used for integrating the microspine array in a consistent, standardized manner. Half of the compliant mechanism contains holes that mechanically anchor it into the silicone limb, highlighted in Fig. 3b), preventing it from being freely pulled out of the limb during microspine gripping. This anchoring method is essential for ensuring the microspine does not come loose over time. The remaining, exposed half of the mechanism contains the microspine.
@@ II-D Soft Robot Design
A tetherless, three limb MTA SoRo with on-board power and processing with AprilTags on each limb is used as the experimental prototype. The components of the physical robot are shown in Fig. 6. Outward trapezoid cavities are introduced on the underside of each limb to provide optimal stiffness and curling ability. This allows the robot to lift the limb and electronic payload. The use of MTA for body deformation enables reliable and efficient limb actuation. The reader may refer to [32] for more details about the topology and morphology design of such modular reconfigurable soft robots.
@@ III EXPERIMENTATION
III-A Experimental Configurations
The baseline experimental configuration with Zero Microspine Limbs, 0ML, is a three-limb MTA SoRo used in [33]. This is compared against two other configurations. The first is One Microspine Limb, 1ML, with the microspine array affixed to a single limb. The second is Two Microspine Limb, 2ML, with microspine arrays equipped on two limbs. In all these configurations, the microspines are angled towards the surface opposite the expected direction of movement. The CAD models of the three prototypes are shown in Fig. 7.
@@ III-A Experimental Configurations
A push-pull translation gait, Fig.  8, is used for the experiments where the gait sequence involves actuation of one limb followed by actuation of the two relaxed limbs. For this locomotion gait, both 0ML and 1ML maintain the same starting pose for a desired path as the single limb that is first actuated in each robot aligns with the intended direction of movement. However, the starting orientation for 2ML is flipped by $180\degree$ so that the two limbs that actuate together face forward. In this configuration, the center point equidistant from the tips of the two microspine limbs faces the intended direction of movement instead of a single limb. This gait was used on all surfaces for all the three robot configurations to facilitate fair comparison. However, an improved performance is possible with an optimized gait that can be synthesized for each robot configuration and surface combination. This may be explored through gait discovery described in [33].
Figure 8: Translation gait: one limb (in maroon) actuates while the other two relax, then the previously relaxed limbs actuate/curl while the first relaxes.
III-B Experimental Setup and Field Experiments
The experiments are carried out with an overhead camera attached to a tripod next to the test area, shown in Fig. 9. The starting pose, position and orientation, is same for each SoRo to ensure consistency. Tracking is performed with an AprilTag fixed to the end of each limb serving as fiducial markers. The average displacement per gait as well as overall displacement for each trial run is recorded.
Figure 9: Example of experimental setup for outside field locations in a forest with a dirt patch. The overhead camera is orthogonal to the ground and approximately 4’ above the surface.
The field experiments are performed on four surfaces of increasing roughness and unstructured nature - uniform concrete, partially uniform brick, granular compact sand with pebbles, and a non-uniform forest floor containing leaf litter and large tree roots. Three trials (60 gaits per trial) are performed for each configuration and surface, resulting in a total of 36 trials. A 36h11 family AprilTag from the AprilTag visual fiducial system [34] is attached to each limb to aid in tracking. The tracking algorithm and data processing code is available at github.com/AgileRoboticsLab/SoftRobotics-Microspines.